OpenStudy (anonymous):

can someone help this nerdy girl? what are the solutions of (x+3)^2=13

OpenStudy (anonymous):

(x+3)^2=13 x+3 = -sqrt13 or sqrt13 x=-3-sqrtr13 or -3+sqrt13

OpenStudy (anonymous):

can u explain the last step please?

OpenStudy (anonymous):

x+3 = -sqrt13 or x+3 = sqrt13 x+3-3=-3-sqrt13 or x+3-3=-3+sqrt13 x=-3-sqrtr13 or -3+sqrt13